# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ 3.2 Slab-experts
$\leq-\sum\limits_{x\in K}\log(\sigma(y_{x}\langle x,{\theta^{\star}}\rangle))-\log\left(\int\limits_{\mathbb{R}^{d}}\exp\left(\sum\limits_{x\in K}\frac{1}{1+\exp(y_{x}\langle x,{\theta^{\star}}\rangle)}\cdot(y_{x}\langle x,\theta-{\theta^{\star}}\rangle)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{1}{8}\cdot(\theta-{\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right)(\theta-{\theta^{\star}})\right)\mu_{K}(d\theta)\right).$
Using Lemma 3.3 to compute the integral of the quadratic form we have
$-\log\left(\int\limits_{\mathbb{R}^{d}}\exp\left(\sum\limits_{x\in K}\frac{1}{1+\exp(y_{x}\langle x,{\theta^{\star}}\rangle)}\cdot(y_{x}\langle x,\theta-{\theta^{\star}}\rangle)-\frac{1}{8}\cdot(\theta-{\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right)(\theta-{\theta^{\star}})\right)\mu_{K}(d\theta)\right)$
$=-\log\left(\left(\frac{\alpha}{\pi}\right)^{\frac{d}{2}}\cdot\left(\frac{\pi}{\alpha+\frac{1}{8}}\right)^{\frac{d}{2}}\exp\left(-\min\limits_{\theta\in\mathbb{R}^{d}}\left(\alpha\theta^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right)\theta+\frac{1}{8}(\theta-{\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right)(\theta-{\theta^{\star}})\right.\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.\left.-\sum\limits_{x\in K}\frac{1}{1+\exp(y_{x}\langle x,{\theta^{\star}}\rangle)}\cdot(y_{x}\langle x,\theta-{\theta^{\star}}\rangle)\right)\right)\right)$
@@
Proof of Theorem 3.2 We first upper bound the total loss by the loss of slab-expert $B_{{\theta^{\star}},\alpha}$ by the definition of the exponential weights algorithm,
$\begin{split}-\sum\limits_{t=1}^{T}\log\tilde{\rho}_{t}(x_{t},y_{t})&=-\sum\limits_{t=1}^{T}\log\left(\frac{\sum\limits_{B\in\mathcal{A}}\exp\left(\sum\limits_{i=1}^{t}\log B(x_{i},y_{i})\right)}{\sum\limits_{B\in\mathcal{A}}\exp\left(\sum\limits_{i=1}^{t-1}\log B(x_{i},y_{i})\right)}\right)\\ &=-\log\left(\frac{1}{|\mathcal{A}|}\sum\limits_{B\in\mathcal{A}}\exp\left(\sum\limits_{i=1}^{T}\log B(x_{i},y_{i})\right)\right)\\ &\leq-\log\left(\exp\left(\sum\limits_{i=1}^{T}\log B_{{\theta^{\star}},\alpha}(x_{i},y_{i})\right)\right)+\log|\mathcal{A}|\\ &\leq-\sum\limits_{i=1}^{T}\log B_{{\theta^{\star}},\alpha}(x_{i},y_{i})+9d\log(eT/d),\end{split}$ (24)
where the last inequality is due to Lemma 3.2. For the slab-expert $B_{{\theta^{\star}},\alpha}$ with $\alpha=1/T^{3}$, it is clear that inside the slab $K_{\theta^{\star}}$ as defined in Definition 3.2, the assumption of Proposition 3.1 is satisfied, thus
@@
$-\sum\limits_{t=1}^{T}\log\tilde{\rho}_{t}(x_{t},y_{t})\leq-\sum\limits_{t=1}^{T}\log(\sigma(y_{t}\langle x_{t},\theta^{\star}\rangle))+11d\log(eT).$
@@ 4.1 Regret bounds for hinge loss
In this section, we first obtain new margin-dependent bounds for the hinge loss in the sequential learning setup using an exponential weight algorithm; then, we proceed to use a transductive prior to achieving an assumption-free regret bound. In the classification setup with classes $1,-1$, given $\gamma>0$, we define the normalized hinge loss of a predictor $\widehat{f}$ as follows
@@
$\mu(\theta)=\sqrt{\det(\lambda A)}\cdot\exp\left(-\pi\lambda\theta^{\top}A\theta\right).$ (29)
@@
$\widehat{f}(x_{t})=\mathbb{E}_{\theta\sim\rho_{t}}\operatorname{clip_{\gamma}}\left(\langle x_{t},\theta\rangle\right).$ (30)
The computation of such prediction depends on the ability of sampling from $\rho_{t}$, which is a log-concave measure. Indeed, once we can approximately sample from $\rho_{t}$, we can use a simple estimate based on Hoeffding’s inequality to identify the number of samples required from $\rho_{t}$ in order to approximate $\widehat{f}(x_{t})$ with the desired accuracy.
Proposition 4.1.
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}(x_{t}))_{+}}{\gamma}$ $\leq(1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\pi\gamma\sqrt{\beta}}\sum\limits_{t=1}^{T}\|A^{-1/2}x_{t}\|_{2}$
$\qquad+\frac{\pi\lambda(\theta^{\star})^{\top}A\theta}{\eta\gamma}+\frac{d}{2\eta\gamma}\left(\log\left(\frac{\beta}{\lambda}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right)\Bigg{)}.$
@@
$\sum\limits_{t=1}^{T}(\gamma-y_{t}\widehat{f}(x_{t}))_{+}\leq\sum\limits_{t=1}^{T}\mathbb{E}_{\theta\sim\rho_{t}}(\gamma-y_{t}\operatorname{clip_{\gamma}}\left(\langle x_{t},\theta\rangle\right))_{+}.$ (31)
Using Bernstein’s inequality for the moment generating function and $\mathbb{E}_{\theta\sim\rho}(\gamma-\operatorname{clip_{\gamma}}\left(y\langle x,\theta\rangle\right))_{+}^{2}\leq 2\gamma\mathbb{E}_{\theta\sim\rho}(\gamma-\operatorname{clip_{\gamma}}\left(y\langle x,\theta\rangle\right))_{+}$, we have for any distribution $\rho$ over $\mathbb{R}^{d}$,
@@
Combining this inequality with (31) and using Lemma 2.1, we obtain
$\sum\limits_{t=1}^{T}\left(1-\frac{\eta\gamma}{1-2\eta\gamma/3}\right)\cdot(\gamma-y_{t}\widehat{f}(x_{t}))_{+}$ $\leq-\sum\limits_{t=1}^{T}\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\exp\left(-\eta(\gamma-\operatorname{clip_{\gamma}}\left(y_{t}\langle x_{t},\theta\rangle\right))_{+}\right)\right)$
@@
By Lemma 2.2, we have
$-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\eta\sum\limits_{t=1}^{T}(\gamma-y_{t}\langle x_{t},\theta\rangle)_{+}\right)\right)\leq\mathbb{E}_{\theta\sim\mu^{\star}}\sum\limits_{t=1}^{T}(\gamma-y_{t}\langle x_{t},\theta\rangle)_{+}+\frac{\mathcal{KL}(\mu^{\star}||\mu)}{\eta}.$
@@
$\mathcal{KL}(\mu^{\star}||\mu)=\pi\lambda(\theta^{\star})^{\top}A\theta+\frac{d}{2}\left(\log\left(\frac{\beta}{\gamma}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right).$
@@
$\sum\limits_{t=1}^{T}\left(1-\frac{\eta\gamma}{1-2\eta\gamma/3}\right)\cdot\frac{(\gamma-y_{t}\widehat{f}(x_{t}))_{+}}{\gamma}$ $\leq\sum\limits_{t=1}^{T}\frac{(1-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\pi\gamma\sqrt{\beta}}\sum\limits_{t=1}^{T}\|A^{-1/2}x_{t}\|_{2}$
$\qquad+\frac{\pi\lambda(\theta^{\star})^{\top}A\theta}{\eta\gamma}+\frac{d}{2\eta\gamma}\left(\log\left(\frac{\beta}{\lambda}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right).$ (32)
@@
There are some immediate corollaries of the result of Proposition 4.1. First, we consider the standard bounded online setup where the vectors $x_{1},\ldots,x_{T}$ are not known in advance.
Corollary 4.1.
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}(x_{t}))_{+}}{\gamma}\leq(1+2\eta\gamma)\left(\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}}{\gamma}+\frac{\frac{1}{\pi}+\pi+\frac{d}{2}\log\left(1+\eta^{2}T^{2}r^{2}b^{2}\right)}{\eta\gamma}\right).$
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}(x_{t}))_{+}}{\gamma}\leq\frac{7}{3}\left(\frac{1}{\pi}+\pi+\frac{d}{2}\log\left(1+\frac{9T^{2}r^{2}b^{2}}{100\gamma^{2}}\right)\right).$ (33)
Proof. The first claim follows from Proposition 4.1 if we plug in $A=I_{d}$, $\lambda=\frac{1}{b^{2}}$, $\beta=\frac{1}{b^{2}}+\eta^{2}T^{2}r^{2}$. The second part of the claim follows from the fact that in the linearly seperable case with margin $\gamma$ we have $\sum\nolimits_{t=1}^{T}(\gamma-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}=0$. The claim follows. ∎
Finally, we present our main result for the transductive setup, where we employ the transductive prior by selecting a data-dependent matrix $A$ in (29). Additionally, we note that a similar sparsity regret bound can be provided, analogous to the one in Theorem 2.2.
@@ 4.2 Slab-experts for hinge loss
Proof of Lemma 4.1 For any $t\in[T]$, suppose $y_{t}\langle x_{t},{\theta^{\star}}\rangle\leq 0$. Since ${\theta^{\star}}$ minimizes the loss, we have
@@
By reorganizing, we obtain the desired result. ∎
Definition 4.1 (Slab-expert for hinge loss).
A slab-expert for the hinge loss $B_{\theta,\alpha,\gamma}$ is an expert that is specified by a vector $\theta\in\mathbb{R}^{d}$, a regularization parameter $\alpha$, and a margin $\gamma>0$. The expert constructs a subset of covariates that belong to a single slab, namely $K_{\theta,\gamma}=\{x_{t}:\left\lvert\langle x_{t},\theta\rangle\right\rvert\leq\gamma T\}$ and predicts labels in the following way: At time step $t$, if $x_{t}\in K_{\theta,\gamma}$, the expert predicts according to $\operatorname{clip_{\gamma}}\left(\mathrm{EWA}_{K_{\theta,\gamma},\alpha}\right)$ with hinge loss and learning rate $\eta$ on $x_{t}$, where $\mathrm{EWA}_{K_{\theta,\gamma},\alpha}$ is the partial exponential weight expert defined as in Definition 3.1 with subset $K_{\theta,\gamma}$ and regularization parameter $\alpha$. If $x_{t}\notin K_{\theta,\gamma}$, then the expert predicts $\hat{y}_{t}=\gamma\operatorname{sign}(\langle x_{t},\theta\rangle)$. We denote by $B_{\theta,\alpha,\gamma}(x_{t})$ the value predicted on receiving the covariate $x_{t}$.
@@
[…]
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}$ $\leq(1+2\eta\gamma)^{2}\left(\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},{\theta^{\star}}\rangle)_{+}}{\gamma}\right)+\frac{12d}{\eta\gamma}(1+2\eta\gamma)^{2}\log(\pi T)$
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}\leq 10+31d\log(\pi T).$ (35)
@@
Proof of Theorem 4.1 Similar to the logistic case, we first upper bound the total loss by the loss of slab-expert $B_{{\theta^{\star}},\alpha,\gamma}$ as defined in Definition 4.1. First, observe that Jensen’s inequality implies
@@
Combining this inequality with (36) and using Lemma Lemma 2.1, we obtain
$\sum\limits_{t=1}^{T}\left(1-\frac{\eta\gamma}{1-2\eta\gamma/3}\right)\cdot(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}$ $\leq-\sum\limits_{t=1}^{T}\frac{1}{\eta}\log\left(\mathbb{E}_{A\sim p_{t}}\exp\left(-\eta(\gamma-y_{t}B(x_{t}))_{+}\right)\right)$
@@
$\leq\sum\limits_{i=1}^{T}(\gamma-y_{i}B_{{\theta^{\star}},\alpha,\gamma}(x_{i}))_{+}+\frac{9d}{\eta}\log(eT/d),$
where the last inequality is due to Lemma 3.2. We apply Proposition 4.1 to the slab-expert $B_{{\theta^{\star}},\alpha,\gamma}$ with $\alpha=\pi\lambda=1/(\gamma^{2}T^{3})$ and $A=\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}xx^{\top}$ to obtain
$\begin{split}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}B_{{\theta^{\star}},\alpha,\gamma}(x))_{+}}{\gamma}&\leq(1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}\langle x,\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\pi\gamma\sqrt{\beta}}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\|A^{-1/2}x\|_{2}\\ &\qquad+\frac{\pi\lambda(\theta^{\star})^{\top}A{\theta^{\star}}}{\eta\gamma}+\frac{d}{2\eta\gamma}\left(\log\left(\frac{\beta}{\lambda}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right)\Bigg{)}.\end{split}$ (37)
@@
$\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}B_{{\theta^{\star}},\alpha,\gamma}(x))_{+}}{\gamma}$ $\leq(1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}\langle x,\theta^{\star}\rangle)_{+}}{\gamma}+\sqrt{d}+\frac{1}{\eta\gamma}+\frac{2d}{\eta\gamma}\log\left(\pi T\right)\Bigg{)}$
@@
Meanwhile, for all the covariates outside the slab, the prediction of $B_{{\theta^{\star}},\alpha}$ coincides with the linear separator with vector ${\theta^{\star}}$. Moreover, $y_{x}\langle x,{\theta^{\star}}\rangle>0$ for all $x\in\{x_{t}\}_{t\in[T]}\setminus K_{\theta^{\star}}$. Thus, we have as a result
$-\sum_{x\in\{x_{t}\}_{t\in[T]}\setminus K_{\theta^{\star}}}\frac{(\gamma-y_{x}B_{{\theta^{\star}},\alpha,\gamma}(x))_{+}}{\gamma}=0.$ (38)

### v3

@@ 3.2 Slab-experts
[…]
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{1}{8}\cdot(\theta-{\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right)(\theta-{\theta^{\star}})\right)\mu_{K}(d\theta)\right).$
Using Lemma 16 to compute the integral of the quadratic form we have
[…]
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-\frac{1}{8}\cdot(\theta-{\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right)(\theta-{\theta^{\star}})\right)\mu_{K}(d\theta)\right)$
[…]
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.\left.-\sum\limits_{x\in K}\frac{1}{1+\exp(y_{x}\langle x,{\theta^{\star}}\rangle)}\cdot(y_{x}\langle x,\theta-{\theta^{\star}}\rangle)\right)\right)\right)$
@@
Proof of Theorem 13 We first upper bound the total loss by the loss of slab-expert $B_{{\theta^{\star}},\alpha}$ by the definition of the exponential weights algorithm,
$\begin{split}-\sum\limits_{t=1}^{T}\log\tilde{\rho}_{t}(x_{t},y_{t})&=-\sum\limits_{t=1}^{T}\log\left(\frac{\sum\limits_{B\in\mathcal{A}}\exp\left(\sum\limits_{i=1}^{t}\log B(x_{i},y_{i})\right)}{\sum\limits_{B\in\mathcal{A}}\exp\left(\sum\limits_{i=1}^{t-1}\log B(x_{i},y_{i})\right)}\right)\\ &=-\log\left(\frac{1}{|\mathcal{A}|}\sum\limits_{B\in\mathcal{A}}\exp\left(\sum\limits_{i=1}^{T}\log B(x_{i},y_{i})\right)\right)\\ &\leq-\log\left(\exp\left(\sum\limits_{i=1}^{T}\log B_{{\theta^{\star}},\alpha}(x_{i},y_{i})\right)\right)+\log|\mathcal{A}|\\ &\leq-\sum\limits_{i=1}^{T}\log B_{{\theta^{\star}},\alpha}(x_{i},y_{i})+4d\log(eT/d),\end{split}$ (24)
where the last inequality is due to Lemma 14. For the slab-expert $B_{{\theta^{\star}},\alpha}$ with $\alpha=1/T^{3}$, it is clear that inside the slab $K_{\theta^{\star}}$ as defined in Definition 12, the assumption of Proposition 15 is satisfied, thus
@@
$-\sum\limits_{t=1}^{T}\log\tilde{\rho}_{t}(x_{t},y_{t})\leq-\sum\limits_{t=1}^{T}\log(\sigma(y_{t}\langle x_{t},\theta^{\star}\rangle))+6d\log(eT).$
@@ 4.1 Regret bounds for hinge loss
In this section, we first obtain new margin-dependent bounds for the hinge loss in the sequential learning setup using an exponential weight algorithm; then, we proceed to use a transductive prior to achieve an assumption-free regret bound. In the classification setup with classes 1,11,-1, given $\gamma>0$, we define the normalized hinge loss of a predictor $\widehat{f}$ as follows
@@
$\mu(\theta)=\sqrt{\frac{\det(\lambda A)}{\pi^{d}}}\cdot\exp\left(-\pi\lambda\theta^{\top}A\theta\right).$ (29)
@@
$\widehat{f}_{t}(x_{t})=\mathbb{E}_{\theta\sim\rho_{t}}\operatorname{clip_{\gamma}}\left(\langle x_{t},\theta\rangle\right).$ (30)
The computation of such prediction depends on the ability of sampling from $\rho_{t}$, which is a log-concave measure. Indeed, once we can approximately sample from $\rho_{t}$, we can use a simple estimate based on Hoeffding’s inequality to identify the number of samples required from $\rho_{t}$ in order to approximate $\widehat{f}_{t}(x_{t})$ with the desired accuracy.
Proposition 17.
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}$ $\leq(1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\pi\gamma\sqrt{\beta}}\sum\limits_{t=1}^{T}\|A^{-1/2}x_{t}\|_{2}$
$\qquad+\frac{\pi\lambda(\theta^{\star})^{\top}A\theta^{\star}}{\eta\gamma}+\frac{d}{2\eta\gamma}\left(\log\left(\frac{\beta}{\lambda}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right)\Bigg{)}.$
@@
$\sum\limits_{t=1}^{T}(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}\leq\sum\limits_{t=1}^{T}\mathbb{E}_{\theta\sim\rho_{t}}(\gamma-y_{t}\operatorname{clip_{\gamma}}\left(\langle x_{t},\theta\rangle\right))_{+}.$ (31)
Using Bernstein’s inequality for the moment generating function and
$\mathbb{E}_{\theta\sim\rho}(\gamma-\operatorname{clip_{\gamma}}\left(y\langle x,\theta\rangle\right))_{+}^{2}\leq 2\gamma\mathbb{E}_{\theta\sim\rho}(\gamma-\operatorname{clip_{\gamma}}\left(y\langle x,\theta\rangle\right))_{+},$
we have for any distribution $\rho$ over $\mathbb{R}^{d}$,
@@
Combining this inequality with (31) and using Lemma 4, we obtain
$\sum\limits_{t=1}^{T}\left(1-\frac{\eta\gamma}{1-2\eta\gamma/3}\right)(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}$ $\leq-\sum\limits_{t=1}^{T}\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\exp\left(-\eta(\gamma-\operatorname{clip_{\gamma}}\left(y_{t}\langle x_{t},\theta\rangle\right))_{+}\right)\right)$
@@
By Lemma 5, we have
$\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\eta\sum\limits_{t=1}^{T}(\gamma-y_{t}\langle x_{t},\theta\rangle)_{+}\right)\right)\leq\mathbb{E}_{\theta\sim\mu^{\star}}\sum\limits_{t=1}^{T}(\gamma-y_{t}\langle x_{t},\theta\rangle)_{+}+\frac{\mathcal{KL}(\mu^{\star}||\mu)}{\eta}.$
@@
$\mathcal{KL}(\mu^{\star}||\mu)=\pi\lambda(\theta^{\star})^{\top}A\theta^{\star}+\frac{d}{2}\left(\log\left(\frac{\beta}{\gamma}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right).$
@@
$\hskip-72.26999pt\sum\limits_{t=1}^{T}\left(1-\frac{\eta\gamma}{1-2\eta\gamma/3}\right)\cdot\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}$
$\leq\sum\limits_{t=1}^{T}\frac{(1-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\pi\gamma\sqrt{\beta}}\sum\limits_{t=1}^{T}\|A^{-1/2}x_{t}\|_{2}+\frac{\pi\lambda(\theta^{\star})^{\top}A\theta^{\star}}{\eta\gamma}$
$\quad+\frac{d}{2\eta\gamma}\left(\log\left(\frac{\beta}{\lambda}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right).$ (32)
@@
There are some immediate corollaries of the result of Proposition 17. First, we consider the standard bounded online setup where the vectors $x_{1},\ldots,x_{T}$ are not known in advance.
Corollary 18.
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}\leq(1+2\eta\gamma)\left(\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}}{\gamma}+\frac{\frac{1}{\pi}+\pi+\frac{d}{2}\log\left(1+\eta^{2}T^{2}r^{2}b^{2}\right)}{\eta\gamma}\right).$
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}\leq\frac{7}{3}\left(\frac{1}{\pi}+\pi+\frac{d}{2}\log\left(1+\frac{9T^{2}r^{2}b^{2}}{100\gamma^{2}}\right)\right).$ (33)
Proof. The first claim follows from Proposition 17 if we plug in $A=I_{d}$, $\lambda=\frac{1}{b^{2}}$, $\beta=\frac{1}{b^{2}}+\eta^{2}T^{2}r^{2}$. The second part of the claim follows from the fact that in the linearly separable case with margin $\gamma$ we have $\sum\nolimits_{t=1}^{T}(\gamma-y_{t}\langle x_{t},\theta^{\star}\rangle)_{+}=0$. The claim follows. ∎
Finally, we present our main result for the transductive setup, where we employ the transductive prior by selecting a data-dependent matrix $A$ in (29). Additionally, we note that a similar sparsity regret bound can be provided, analogous to the one in Theorem 3.
@@ 4.2 Slab-experts for hinge loss
Proof of Lemma 19 For any $t\in[T]$, suppose $y_{t}\langle x_{t},{\theta^{\star}}\rangle\leq 0$. Since $\theta^{\star}$} minimizes the loss, we have
@@
Reorganizing terms, we obtain the desired result. ∎
Definition 20 (Slab-expert for hinge loss).
A slab-expert for the hinge loss $B_{\theta,\alpha,\gamma}$ is an expert that is specified by a vector $\theta\in\mathbb{R}^{d}$, a regularization parameter $\alpha$, and a margin $\gamma>0$. The expert constructs a subset of covariates that belong to a single slab, namely $K_{\theta,\gamma}=\{x_{t}:\left\lvert\langle x_{t},\theta\rangle\right\rvert\leq\gamma T\}$ and predicts labels in the following way: At time step $t$, if $x_{t}\in K_{\theta,\gamma}$, the expert predicts according to $\operatorname{clip_{\gamma}}\left(\mathrm{EWA}_{K_{\theta,\gamma},\alpha}\right)$ with hinge loss and learning rate $\eta$ on $x_{t}$, where $\mathrm{EWA}_{K_{\theta,\gamma},\alpha}$ is the partial exponential weight expert defined as in Definition 11 with subset $K_{\theta,\gamma}$ and regularization parameter $\alpha$. If $x_{t}\notin K_{\theta,\gamma}$, then the expert predicts $\hat{y}_{t}=\gamma\operatorname{sign}(\langle x_{t},\theta\rangle)$. We denote by $B_{\theta,\alpha,\gamma}(x_{t})$ the value predicted on receiving the covariate $x_{t}$.
@@
   The learner predicts with aggregation on the slab-experts $\widehat{f}_{t}(x_{t})=\sum\limits_{B\in\mathcal{A}}p_{t}(B)B(x_{t})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ }p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}-\eta(\gamma-y_{i}B(x_{i}))_{+}\right).$
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}$ $\leq(1+2\eta\gamma)^{2}\left(\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\langle x_{t},{\theta^{\star}}\rangle)_{+}}{\gamma}\right)+\frac{7d}{\eta\gamma}(1+2\eta\gamma)^{2}\log(\pi T)$
@@
$\sum\limits_{t=1}^{T}\frac{(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}}{\gamma}\leq 9+60d\log(\pi T).$ (35)
@@
Proof of Theorem 21 Similar to the logistic case, we first upper bound the total loss by the loss of slab-expert $B_{{\theta^{\star}},\alpha,\gamma}$ as defined in Definition 20. First, observe that Jensen’s inequality implies
@@
Combining this inequality with (36) and using Lemma Lemma 4, we obtain
$\hskip-72.26999pt\sum\limits_{t=1}^{T}\left(1-\frac{\eta\gamma}{1-2\eta\gamma/3}\right)\cdot(\gamma-y_{t}\widehat{f}_{t}(x_{t}))_{+}$
$\leq-\sum\limits_{t=1}^{T}\frac{1}{\eta}\log\left(\mathbb{E}_{A\sim p_{t}}\exp\left(-\eta(\gamma-y_{t}B(x_{t}))_{+}\right)\right)$
@@
$\leq\sum\limits_{i=1}^{T}(\gamma-y_{i}B_{{\theta^{\star}},\alpha,\gamma}(x_{i}))_{+}+\frac{4d}{\eta}\log(eT/d),$
where the last inequality is due to Lemma 14. We apply Proposition 17 to the slab-expert $B_{{\theta^{\star}},\alpha,\gamma}$ with $\alpha=\pi\lambda=1/(\gamma^{2}T^{3})$ and $A=\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}xx^{\top}$ to obtain
$\begin{split}\hskip 72.26999pt&\hskip-72.26999pt\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}B_{{\theta^{\star}},\alpha,\gamma}(x))_{+}}{\gamma}\\ &\leq(1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}\langle x,\theta^{\star}\rangle)_{+}}{\gamma}+\frac{1}{\pi\gamma\sqrt{\beta}}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\|A^{-1/2}x\|_{2}\\ &\qquad+\frac{\pi\lambda(\theta^{\star})^{\top}A{\theta^{\star}}}{\eta\gamma}+\frac{d}{2\eta\gamma}\left(\log\left(\frac{\beta}{\lambda}\right)-\frac{1}{2}+\frac{\lambda}{2\beta}\right)\Bigg{)}.\end{split}$ (37)
@@
$\hskip-72.26999pt\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}B_{{\theta^{\star}},\alpha,\gamma}(x))_{+}}{\gamma}$
$\leq(1+2\eta\gamma)\cdot\Bigg{(}\sum\limits_{x\in K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}\langle x,\theta^{\star}\rangle)_{+}}{\gamma}+\sqrt{d}+\frac{1}{\eta\gamma}+\frac{2d}{\eta\gamma}\log\left(\pi T\right)\Bigg{)}$
@@
Meanwhile, for all the covariates outside the slab, the prediction of $B_{{\theta^{\star}},\alpha}$ coincides with the linear separator with vector $\theta^{\star}$}. Moreover, $y_{x}\langle x,{\theta^{\star}}\rangle>0$ for all $x\in\{x_{t}\}_{t\in[T]}\setminus K_{{\theta^{\star}},\gamma}$. Thus, we have as a result
$\sum_{x\in\{x_{t}\}_{t\in[T]}\setminus K_{{\theta^{\star}},\gamma}}\frac{(\gamma-y_{x}B_{{\theta^{\star}},\alpha,\gamma}(x))_{+}}{\gamma}=0.$ (38)